callielovesyhuu:

Greg earned $900 for washing 30 windows last month. If Greg's wage is increased by $2 per window this month, how much will he earn for washing 30 windows this month?

callielovesyhuu:

@Warriorz13 @wo1f0mon @Allison @SkyVoltage43 @theDeviliscoming

theDeviliscoming:

900/30=30 30+2=32 32*30=960 Greg will earn $960 for washing 30 windows

ldenault:

he will earn $960

HuskyNation:

$960